OpenStudy (anonymous):

LAST 2 PLEASE HELP, simplify the product using a table. 8. (5h + 4)(3h + 6) __________ 3h | 6 __________ 5h _________ 4 A. 15h2 + 42h + 24 B. 15h2 – 42h + 24 C. 15h2 + 18h – 24 D. 15h2 – 18h – 24 9. (–2h + 4)(5h + 7) _______ 5h 7 _______ -2h ________ 4 A. –10h2 + 34h – 28 B. –10h2 + 6h + 28 C. –10h2 – 6h + 28 D. –10h2 – 34h – 28

OpenStudy (anonymous):

i think 8 is D and 9 is D?

OpenStudy (whpalmer4):

\[(5h+4)(3h+6) = 5h(3h+6) + 4(3h+6) \]by distributive property of multiplication \[15h^2+30h+12h+24 = 15h^2+42h+24\]

OpenStudy (whpalmer4):

\[(-2h+4)(5h+7) = -2h(5h+7)+4(5h+7)\]Using the same approach, what does that give you?

OpenStudy (anonymous):

// i don't understand why you need a table, but this is how i did it. (5h + 4)(3h + 6) (5h*3h)+(4*3h)+(5h*6)+(4*6) 15h^2+12h+30h+24 15h^2+42h+24

OpenStudy (anonymous):

thank you guys! help with 9?

OpenStudy (whpalmer4):

Yeah, I was a bit puzzled by the table approach, too. Maybe if we saw the text it would be clearer...

OpenStudy (whpalmer4):

I showed you what to do with 9...continue the work and tell us what you get...

OpenStudy (anonymous):

@whpalmer4 you both commented at the same time i didmt realize yours was for #9, thank you!

OpenStudy (whpalmer4):

Okay, now you know, what's the answer? :-)

OpenStudy (anonymous):

is it b/:

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

If you multiply it out correctly, you won't have to ask. Multiply it out and tell me what you get.

OpenStudy (whpalmer4):

Actually writing it out is productive...

OpenStudy (anonymous):

c(:

OpenStudy (whpalmer4):

Is that an emoticon, or an answer choice? \[(-2h+4)(5h+7) = -2h(5h+7)+4(5h+7) = \] \[-2h(5h) -2h(7)+4(5h) + 4(7) = -10h^2-14h+20h+28 =\]\[-10h^2+(20-14)h+28=-10h^2+6h+28\]

OpenStudy (anonymous):

c is my answer lol!

OpenStudy (whpalmer4):

Can you point out the mistake I made, then? :-)

OpenStudy (anonymous):

i don't think you did. i suck at math so im probably wrong!